OpenStudy (anonymous):

the inverse of matrix 2x2 [3 2] [3 5]

OpenStudy (anonymous):

help

OpenStudy (freckles):

if the determinant exist then the inverse matrix exist but the inverse of a 2 by 2 matrix where the determinant exist is \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]^{-1}=\frac{1}{ad-bc} \left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]